# Edge Learning for 6G-enabled Internet of Things: A Comprehensive Survey of Vulnerabilities, Datasets, and Defenses

### v1

@@ IX-B2 Datasets Formation/Availability for Edge Learning
In addition to introducing security into AI, other challenges may arise along the way, including the availability and the creation of dependable edge learning datasets. We highlight two important issues, namely managing complexity and ensuring reliable data resource availability.
  • Learning Complexity: In a complex environment such as mobile communication systems, which is likely to be even more complex with the projected introduction of a range of technologies for future networks such as 6G, and given the limited hardware capabilities of IoT devices and cellular entities, particularly in terms of processing power, battery life, and storage, this is going to be a serious problem in terms of how well these entities can handle complexity in the interests of functionality. Explainable AI (XAI) can solve part of the problem, unlike the black-box concept, as it effectively introduces understanding, management, and a level of trust in AI models, making it easier to reduce complexity without affecting accuracy. XAI is an important tool for managing complexity in mobile communication systems and other complex environments. By providing transparency, understanding, and trust in AI models, XAI can help stakeholders navigate complex environments while maintaining functionality and accuracy.
  • High-Quality Data Availability: The FL paradigm solves the problem of compromising the privacy of private data in the learning phase. However, as we have seen in the previous sections, this paradigm is subject to different types of threats. One of them is related to data, which can be a significant danger for future wireless networks since distributed learning is supposed to be a key component of future networks. It can be tricky to trust models trained on data that the operator has not seen or validated, such as when relying on UEs for the training task. Similarly, even when dealing with internal components such as gNBs and Intelligent Radios (IRs), it can be dangerous to trust data collected from the surrounding environment and used for training directly without the proper validations. Techniques that introduce a confidence level and validations must be applied before incorporating the models directly into the networks. Overall, the availability of high-quality data and the use of techniques that introduce a confidence level and validations are critical to ensuring the success of federated learning in wireless networks. By adopting these techniques, operators can trust the models generated by the learning process and use them to make informed decisions about the network’s configuration and performance.
@@ IX-B3 ML Vulnerabilities Elimination and Defense Strategies Implementations
It is difficult to develop a security solution that guarantees equal levels for conflicting aspects. Here we discuss the trade-offs that can arise when trying to eliminate ML vulnerabilities and develop security solutions.
@@
  • Security vs. Performance and/or Profit: The very high data rates promised by 6G will enable a variety of time-sensitive IoE applications, requiring real-time security techniques to be deployed in turn, as conventional security schemes are subject to short-term and long-term failure under such circumstances. This results in different trade-offs, on the one hand, between security and performance, e.g., between privacy and accuracy in FL, since introducing a lot of noise to protect the data will result in reduced accuracy, and vice versa. On the other hand, the effort for real-time security techniques is costly and depends on what the telecom industry is willing to pay and what are their priorities. The effort for developing and implementing real-time security techniques is costly, and the telecom industry may have to prioritize security, performance, and profitability. In some cases, the industry may decide to prioritize performance or profitability over security, which could result in compromises in terms of the security of the network and the data transmitted over it. Therefore, it is essential to strike a balance between security, performance, and profitability while deploying real-time security techniques for 6G networks. This will require careful consideration of the specific requirements and constraints of different applications and industries and collaboration between stakeholders in the telecom industry, academia, and government.
@@
  • Simplicity vs. Complexity: Since future networks are expected to be more complex than previous generations, which is natural, and this has persisted in the evolution of all preceding generations. A trivial conclusion that could be drawn directly is that security solutions need to be complex as well, however, this can be seen as a double-edged sword, since, in order to keep the system stable, more secure, and have fewer vulnerabilities, simplicity, and transparency are keys [44]. In the case of 6G-enabled IoT networks, which are expected to be more complex, it is important to consider both simplicity and complexity in security solutions. One approach could be to implement a layered security model, where multiple layers of security are used to protect against different types of attacks. Each layer can be designed with a simple and transparent solution that is easy to understand and audit, while also being integrated into a larger, more complex security architecture. ML-based security is a promising approach for detecting zero-day attacks, which are attacks that exploit vulnerabilities that are unknown to security experts. However, as mentioned, zero-day attacks that specifically target the ML algorithms themselves can be difficult to detect. This highlights the need for ongoing research and development in security solutions to stay ahead of emerging threats.
@@
  • Backwards Compatibility vs. Future Security: In addition to the increased complexity, maintaining backward compatibility is another feature that persists in all previous generations. While it has its benefits, including reduced deployment costs compared to a new standalone deployment and ease of upgrading, the backward compatibility functionality potentially exposes old vulnerabilities. In addition, AI-based security must also take into account some of the previous generation’s parameters in order to function properly. This may be something to be carefully examined in the future. Overall, balancing the need for backward compatibility with the need for future security is a complex challenge that requires careful consideration. It may be necessary to prioritize security over backward compatibility in certain cases, but in other cases, it may be possible to reach a trade-off between the two.
@@ IX-B4 Large language models at the Edge
Large language models, like OpenAI’s GPT-4, FalconLLM, and Bert, are increasingly being used in a wide range of applications, from content generation to customer service and much more [269, 270, 271]. However, deploying large language models at the edge for 6G-enabled IoT systems introduces several challenges. First, security and privacy issues are magnified in a 6G environment due to the sheer volume and variety of devices [272]. The immense data flow, which includes sensitive data, presents a significant risk of cyber threats and vulnerabilities. Second, despite 6G’s promise of higher data rates and lower latencies, managing the scale and complexity of these models can be a challenge due to the inherent resource constraints of IoT devices, potentially affecting model performance and efficiency. The high density of IoT devices in 6G networks could also strain network resources and introduce latency issues, impacting real-time responses and decision-making [95]. Third, maintaining consistency of the model’s deployment across a diverse range of IoT devices is difficult, and software updates become challenging due to the vast distribution of devices. Therefore, the deployment of large language models at the edge of 6G IoT environments necessitates careful attention to these challenges to ensure security, efficiency, and compliance.
@@ X Conclusions
While 5G is being deployed and commercialized around the world, researchers are getting excited about what 6G can and should be. One of the most endorsed views is that AI should be a core component of 6G rather than just an add-on utility. edge learning involves making the edge of the network intelligent, where models are trained at the edge using coordinated distributed learning paradigms such as FL, with data available across a variety of edge devices. However, given the existing vulnerabilities in ML, its adoption in the absence of adequate security considerations can expose the network to various threats. In this paper, we have surveyed the state-of-the-art of existing vulnerabilities and defenses of federated machine learning for 6G-enabled IoTs. We have summarized the existing surveys on machine learning for 6G–IoT security as well as machine learning-associated threats in three different learning modes, namely, centralized, federated, and distributed. Through extensive research and analysis that has been conducted, we have classified the threat models against machine learning into eight categories, including, backdoor attacks, adversarial examples, combined attacks, poisoning attacks, Sybil attacks, byzantine attacks, inference attacks, and dropping attacks. In addition, we have analyzed the state-of-the-art defense methods against federated machine learning vulnerabilities. Finally, as new attacks and defense technologies are realized, new research and future overall prospects for 6G-enabled IoTs are discussed. There still exist several challenging research areas on new attacks and defense technologies, which should be further investigated in the near future.

### v2

@@ IX-B2 Datasets Formation/Availability for Edge Learning
In addition to introducing security into AI, other challenges may arise along the way, including the availability and the creation of dependable Edge Learning datasets. We highlight two important issues: managing complexity and ensuring reliable data resource availability.
  • Learning Complexity: In a complex environment such as mobile communication systems, which is likely to be even more complex with the projected introduction of a range of technologies for future networks such as 6G, and given the limited hardware capabilities of IoT devices and cellular entities, particularly in terms of processing power, battery life, and storage, this is going to be a serious problem in terms of how well these entities can handle complexity in the interests of functionality. Explainable AI (XAI) can solve part of the problem, unlike the black-box concept, as it effectively introduces understanding, management, and trust in AI models, making it easier to reduce complexity without affecting accuracy. XAI is an important tool for managing complexity in mobile communication systems and other environments. By providing transparency, understanding, and trust in AI models, XAI can help stakeholders navigate complex environments while maintaining functionality and accuracy.
  • High-Quality Data Availability: The FL paradigm solves the problem of compromising the privacy of private data in the learning phase. However, as we have seen in the previous sections, this paradigm is subject to different threats. One is related to data, which can be a significant danger for future wireless networks since distributed learning is supposed to be a key component of future networks. It can be tricky to trust models trained on data the operator has not seen or validated, such as when relying on UEs for the training task. Similarly, even when dealing with internal components such as gNBs and Intelligent Radios (IRs), it can be dangerous to trust data collected from the surrounding environment and used for training directly without the proper validations. Techniques that introduce a confidence level and validations must be applied before incorporating the models directly into the networks. Overall, high-quality data availability and techniques that introduce a confidence level and validations are critical to ensuring the success of federated learning in wireless networks. By adopting these techniques, operators can trust the models generated by the learning process and use them to make informed decisions about the network’s configuration and performance.
@@ IX-B3 ML Vulnerabilities Elimination and Defense Strategies Implementations
Developing a security solution that guarantees equal levels for conflicting aspects is challenging. Here we discuss the trade-offs that can arise when trying to eliminate ML vulnerabilities and develop security solutions.
@@
  • Security vs. Performance and/or Profit: The high data rates promised by 6G will enable various time-sensitive IoE applications, requiring real-time security techniques to be deployed, as conventional security schemes are subject to short-term and long-term failure under such circumstances. This results in different trade-offs, on the one hand, between security and performance, e.g., between privacy and accuracy in FL, since introducing a lot of noise to protect the data will result in reduced accuracy, and vice versa. On the other hand, the effort for real-time security techniques is costly and depends on what the telecom industry is willing to pay and their priorities. The effort for developing and implementing real-time security techniques is costly, and the telecom industry may have to prioritize security, performance, and profitability. In some cases, the industry may decide to prioritize performance or profitability over security, which could compromise the security of the network and the data transmitted over it. Therefore, it is essential to balance security, performance, and profitability while deploying real-time security techniques for 6G networks. This will require careful consideration of different applications and industries’ specific requirements and constraints and collaboration between telecom, academia, and government stakeholders.
@@
  • Simplicity vs. Complexity: Since future networks are expected to be more complex than previous generations, which is natural, this has persisted in the evolution of all preceding generations. A trivial conclusion that could be drawn directly is that security solutions need to be complex as well; however, this can be seen as a double-edged sword since, to keep the system stable, more secure, and have fewer vulnerabilities, simplicity, and transparency are keys [58]. In the case of 6G-enabled IoT networks, which are expected to be more complex, it is important to consider both simplicity and complexity in security solutions. One approach could be implementing a layered security model, where multiple layers of security are used to protect against different attacks. Each layer can be designed with a simple and transparent solution that is easy to understand and audit while also being integrated into a larger, more complex security architecture. ML-based security is a promising approach for detecting zero-day attacks, which are attacks that exploit vulnerabilities that are unknown to security experts. However, zero-day attacks targeting ML algorithms can be difficult to detect. This highlights the need for ongoing research and development in security solutions to stay ahead of emerging threats.
@@
  • Backwards Compatibility vs. Future Security: In addition to the increased complexity, maintaining backward compatibility is another feature that persists in all previous generations. While it has benefits, including reduced deployment costs compared to a new standalone deployment and ease of upgrading, the backward compatibility functionality potentially exposes old vulnerabilities. In addition, AI-based security must also consider some of the previous generation’s parameters to function properly. This may be something to be carefully examined in the future. Overall, balancing the need for backward compatibility with the need for future security is a complex challenge that requires careful consideration. It may be necessary to prioritize security over backward compatibility in certain cases, but in other cases, it may be possible to reach a trade-off between the two.
@@ IX-B4 Large language models at the Edge
Large language models, like OpenAI’s GPT-4, FalconLLM, and Bert, are increasingly being used in a wide range of applications, from content generation to customer service and much more [296, 297, 298]. However, deploying large language models at the edge for 6G-enabled IoT systems introduces several challenges. First, security and privacy issues are magnified in a 6G environment due to the sheer volume and variety of devices [132]. The immense data flow, which includes sensitive data, presents a significant risk of cyber threats and vulnerabilities. Second, despite 6G’s promise of higher data rates and lower latencies, managing the scale and complexity of these models can be a challenge due to the inherent resource constraints of IoT devices, potentially affecting model performance and efficiency. The high density of IoT devices in 6G networks could also strain network resources and introduce latency issues, impacting real-time responses and decision-making [64]. Third, maintaining consistency of the model’s deployment across a diverse range of IoT devices is difficult, and software updates become challenging due to the vast distribution of devices. Therefore, deploying large language models at the edge of 6G IoT environments necessitates careful attention to these challenges to ensure security, efficiency, and compliance.
IX-B5 Universal Multi-prompt and Multi-model attacks
Creating adversarial attacks for discrete forms of data such as text is substantially more challenging than for continuous data types like images [299]. This is primarily due to the difficulties in generating adversarial samples through gradient-based techniques. Existing successful strategies for attacking text data typically involve heuristic replacement methods at either the character or word level [300]. Recently, Zou et al. [301] presented a novel method for creating adversarial attacks on large language models (LLMs) to generate objectionable content. This method, unlike previous human-centric or inefficient automated techniques, generates adversarial suffixes through automated greedy and gradient-based search techniques, prompting LLMs to produce undesired responses. These generated prompts were highly transferable across several LLMs, including ChatGPT, Bard, Claude, and others, with a notable susceptibility in GPT-based models. The study’s findings advance the understanding of adversarial attacks against LLMs and highlight the need to improve these systems’ resistance to producing inappropriate content. The challenge is developing effective defenses against adversarial attacks that prompt large language models to generate objectionable content.
@@ X Conclusions
While 5G is being deployed and commercialized around the world, researchers are getting excited about what 6G can and should be. One of the most endorsed views is that AI should be a core component of 6G rather than just an add-on utility. Edge learning involves making the edge of the network intelligent, where models are trained at the edge using coordinated distributed learning paradigms such as FL, with data available across various edge devices. However, given the existing vulnerabilities in ML, its adoption without adequate security considerations can expose the network to various threats. In this paper, we have surveyed the state-of-the-art of existing vulnerabilities and defenses of federated machine learning for 6G-enabled IoTs. We have summarized the existing surveys on machine learning for 6G–IoT security and machine learning-associated threats in three different learning modes: centralized, federated, and distributed. Through extensive research and analysis that has been conducted, we have classified the threat models against machine learning into eight categories, including backdoor attacks, adversarial examples, combined attacks, poisoning attacks, Sybil attacks, byzantine attacks, inference attacks, and dropping attacks. In addition, we have analyzed the state-of-the-art defense methods against federated machine learning vulnerabilities. Finally, as new attacks and defense technologies are realized, new research and future overall prospects for 6G-enabled IoTs are discussed. There still exist several challenging research areas on new attacks and defense technologies, which should be further investigated in the near future.